# Study on quantum thermalization from thermal initial states in a superconducting quantum computer

### v1

@@ Abstract
Quantum thermalization in contemporary quantum devices, in particular quantum computers, has recently attracted significant theoretical interest. Unusual thermalization processes, such as the Quantum Mpemba Effect (QME), have been explored theoretically. However, there is a shortage of experimental results due to the difficulty in preparing thermal states. In this paper, we propose a method to address this challenge. Moreover, we experimentally validate our approach using IBM quantum devices, providing results for unusal relaxation in equidistant quenches as predicted for the IBM qubit. We also assess the formalism introduced for the QME, obtaining results consistent with the theoretical predictions. This demonstration underscores that our method can streamline the investigation of thermal states and thermalization in quantum physics.
@@ I Introduction
In this work, we introduce a method for indirectly studying the physics of initial thermal states in any quantum system, provided that we can prepare the eigenstates of the Hamiltonian. We validate this method using two IBM quantum computers, experimentally investigating the potential occurrence of the Quantum Mpemba Effect (QME) in this system. Furthermore, we examine another unusal thermalization process known as Equidistant Quenches [8, 9], which contrasts the relaxation into equilibrium of a pure state with that of an initially equidistant thermal state.
@@
The paper is structured as follows: In Section 2, we introduce the method and present the theoretical results essential for this work. Section 3 outlines the methodology employed in the experiments. The results obtained from the IBM devices are presented in Section 4, followed by the conclusion in Section 5.
@@ II Theoretical Background
$\rho_{th}(T,t)$ $=\sum_{n,\alpha}p_{n}(T)K_{\alpha}(t,t_{0})\ket{n(t_{0})}\bra{n(t_{0})}K^{\dagger}{\alpha}(t,t_{0})$ (4)
@@
Here, $K_{\alpha}$ represents the Kraus operators, satisfying the normalization condition $K_{\alpha}K^{\dagger}_{\alpha}=\mathbb{I}$. Since $\rho{n}(t)$ are the time-dependent density matrices for each eigenstate, we can individually measure the evolution for each |$\ket{n}$ and then add them with probability $p_{n}(T)$, obtaining the equivalent dynamics for any initial thermal state.
@@
III Methods
In this work, we present two experiments, and a schematic representation of all executed circuits can be seen in Figure 1. The first experiment was conducted on qubit 0 of the 5-qubit ibmq_belem device. We prepared |$\ket{0}$ and |$\ket{1}$ along with two different superposition states, |$\ket{\Psi_{1}}=\frac{2}{\sqrt{5}}\ket{0}+\frac{1}{\sqrt{5}}\ket{1}$ and |$\ket{\Psi_{2}}=\frac{3}{\sqrt{10}}\ket{0}+\frac{1}{\sqrt{10}}\ket{1}$. We took 33 snapshots, each separated by 8 $\mu s$. The choice of this number is due to the limitation imposed by the IBM service, allowing only 100 different circuits in each usage of the quantum computer. We needed 3 circuits for each snapshot to perform Quantum State Tomography (QST) [17] for reconstructing the entire density matrix for both pure states, considering their off-diagonal contributions. QST was implemented using the Qiskit Softwate Development Kit (SDK) [18].
For the second experiment, on qubit 0 of the 5-qubit ibmq_manila computer, we prepared |$\ket{0}$ and |$\ket{1}$. We took 100 snapshots with a separation of 32 $\mu s$. As there were no off-diagonal terms contributing to any of the states in this case, QST was not needed. This allowed us to capture a higher number of snapshots. Diagonal density matrix elements were calculated by dividing the number of counts for each state by the total number of shots, $n_{\text{shots}}=8192$, in all circuits for both experiments.
In both experiments, thermal states were constructed as explained before, by adding the results for the pure initial eigenstates |$\ket{0}$ and |$\ket{1}$ at each time step with the corresponding ponderation given by Equation 3. In Table 1, we present the hardware calibration data provided by IBM at the time of the experiment for ibmq_belem and ibmq_manila.
Figure 1: Scheme for the experiments, the particular gates corresponding to the initialization block varies depending on the state and the free evolution time is changed in each snapshot in order to observe the entire thermalization process.
Table 1: Calibration values at the moment of the experiment.
@@ V Discussion
Through simple experiments, we have indirectly tested numerous results of quantum thermodynamics that would otherwise require a challenging experimental setup. This underscores the potential utility of NISQ quantum computers not only as calculating devices but also as systems for studying and testing Open Quantum Systems physics.
@@ Acknowledgements.
We gratefully acknowledge supports from the Basque Country Government (Grant No. IT1470- 22) and Grant No. PGC2018-101355-B-I00 funded by MCIN/AEI/10.13039/501100011033.

### v2

@@ Abstract
Quantum thermalization in contemporary quantum devices, in particular quantum computers, has recently attracted significant theoretical interest. Unusual thermalization processes, such as the Quantum Mpemba Effect (QME), have been explored theoretically. However, there is a shortage of experimental results due to the difficulty in preparing thermal states. In this paper, we propose a method to indirectly address this challenge. Moreover, we experimentally validate our approach using IBM quantum devices, providing results for unusual relaxation in equidistant quenches as predicted for the IBM qubit. We also assess the formalism introduced for the QME, obtaining results consistent with the theoretical predictions. This demonstration underscores that our method can streamline the investigation of thermal states and thermalization in quantum physics.
@@ I Introduction
In this work, we introduce a method for indirectly studying the physics of initial thermal states in any quantum system, provided that we can prepare the eigenstates of the Hamiltonian. We validate this method using two IBM quantum computers, experimentally testing the formalism for the QME in this system. Furthermore, we examine another unusal thermalization process known as Equidistant Quenches [8, 9], which contrasts the relaxation into equilibrium of a pure state with that of an initially equidistant thermal state.
@@
The paper is structured as follows: In Section 2, we introduce the method and present the theoretical results essential for this work. Section 3 outlines the methodology employed in the demonstrations. The results obtained from the IBM devices are presented in Section 4, followed by the conclusion in Section 5.
@@ II Theoretical Background
$\rho_{th}(T,t)$ $=\sum_{n,\alpha}p_{n}(T)K_{\alpha}(t,t_{0})\ket{n(t_{0})}\bra{n(t_{0})}K^{\dagger}_{\alpha}(t,t_{0})$ (4)
@@
Here, $K_{\alpha}$ represents the Kraus operators, satisfying the normalization condition $K_{\alpha}K^{\dagger}_{\alpha}=\mathbb{I}$. Since $\rho_{n}(t)$ are the time-dependent density matrices for each eigenstate, we can individually measure the evolution for each |$\ket{n}$ and then add them with probability $p_{n}(T)$, obtaining the equivalent dynamics for any initial thermal state.
@@
The ocurrence or absence of the effect in a Markovian system can be predicted from the form of the probability vector in terms of the eigenvectors of the master equation, that is,
$\vec{p}(t)=\vec{p}_{eq}(T_{eq})+a_{2}e^{\lambda_{2}t}\vec{v}_{2}(t)+...+a_{n}e^{\lambda_{n}t}\vec{v}_{n}(t).$ (11)
Then, a system will present the QME [3] if
$\lambda_{2}>\lambda_{3}\quad,\quad|a_{2}^{c}|>|a_{2}^{h}|.$ (12)
In the case of a system described by Equation 5 the effect should not appear since we always have that $\lambda_{3}=0$ and $\lambda_{2}<0$ [8].
III IBMQ Demonstrations
In this work, we present two demonstrations, and a schematic representation of all executed circuits can be seen in Figure 1. The first demonstration was conducted on qubit 0 of the 5-qubit ibmq_belem device. We prepared |$\ket{0}$ and |$\ket{1}$ along with two different superposition states, |$\ket{\Psi_{1}}=\frac{2}{\sqrt{5}}\ket{0}+\frac{1}{\sqrt{5}}\ket{1}$ and |$\ket{\Psi_{2}}=\frac{3}{\sqrt{10}}\ket{0}+\frac{1}{\sqrt{10}}\ket{1}$. We took 33 snapshots, each separated by 8 $\mu s$. The choice of this number is due to the limitation imposed by the IBM service, allowing only 100 different circuits in each usage of the quantum computer. We needed 3 circuits for each snapshot to perform Quantum State Tomography (QST) [17] for reconstructing the entire density matrix for both pure states, considering their off-diagonal contributions. QST was implemented using the Qiskit Softwate Development Kit (SDK) [18].
For the second demonstration, on qubit 0 of the 5-qubit ibmq_manila computer, we prepared |$\ket{0}$ and |$\ket{1}$. We took 100 snapshots with a separation of 32 $\mu s$. As there were no off-diagonal terms contributing to any of the states in this case, QST was not needed. This allowed us to capture a higher number of snapshots. Diagonal density matrix elements were calculated by dividing the number of counts for each state by the total number of shots, $n_{\text{shots}}=8192$, in all circuits for both demonstrations.
In both demonstrations, thermal states were constructed as explained before, by adding the results for the pure initial eigenstates |$\ket{0}$ and |$\ket{1}$ at each time step with the corresponding ponderation given by Equation 3. In Table 1, we present the hardware calibration data provided by IBM at the time of the demonstration for ibmq_belem and ibmq_manila.
Figure 1: Scheme for the demonstrations, the particular gates corresponding to the initialization block varies depending on the state and the free evolution time is changed in each snapshot in order to observe the entire thermalization process.
Table 1: Calibration values at the time of the demonstration.
@@ V Discussion
However, it is important to highlight the limitations and weaknesses of the proposed method. Firstly, despite the formal equivalence between eigenstates and thermal state evolution, thermal states are not actually prepared on the quantum system. Therefore, our method is not applicable for quantum algorithms that utilize thermal states as a computational tool. Additionally, ensuring identical interactions with the environment for each eigenstate evolution is necessary. We have addressed this by running the different circuits as closely in time as possible. Further investigation of this method on different quantum systems with higher complexity is needed.
Through simple demonstrations, we have indirectly tested numerous results of quantum thermodynamics that would otherwise require a challenging experimental setup. This underscores the potential utility of NISQ quantum computers not only as calculating devices but also as systems for studying and testing Open Quantum Systems physics.
@@ Acknowledgements.
We would like to thank Dr Dawei Luo for discussions. We gratefully acknowledge supports from the Basque Country Government (Grant No. IT1470- 22) and Grant No. PGC2018-101355-B-I00 funded by MCIN/AEI/10.13039/501100011033.